Lou.. arlöw, а.di isionär mänägert r. Säge. Company, as än opportunity to manufacture. and sell one of two new products for a five-year period. His annual pay raises are i determined by his divisions return on investment (ROI), which has exceeded 25% each of the last three years. He has computed the cost and revenue estimates for each product as follows ProductAProduct B Initial investment: Cost of equipment (zero salvage value) Annual revenues and costs: Sales revenues Variable expenses Depreciation expense Fixed out-of-pocket operating costs S 370, 000 530,000 S 400, 000 510,000 S 180, 000 250,000 $ 74, 000 106,000 $ 85, 000 72,000 I he companys discount rate is 19% Click here to view Exhibit 128-1 and Exhibit 12B-2, to determine the appropriate discount factor using tables. Required 1. Calculate the payback period for each product. 2. Calculate the net present value for each product. 3. Calculate the internal rate of return for each product. 4. Calculate the project profitability index for each product. 5. Calculate the simple rate of return for each product. 6a. For each measure, identify whether Product A or Product B is preferred 6b. Based on the simple rate of return, Lou Barlow would likely:
Answer #1
1) Project A Project B
Payback Period 2.74 Years 2.82 Years
Working:
Payback period is the time upto which initial cost is recovered back.
Project A Project B
Initial Investment a $   3,70,000.00 $   5,30,000.00
Annual cash flows:
Sales revenues $   4,00,000.00 $   5,10,000.00
Variable expense $ -1,80,000.00 $ -2,50,000.00
Fixed-out-of-pocket operating costs $     -85,000.00 $     -72,000.00
Annual cash flows b $   1,35,000.00 $   1,88,000.00
Payback Period (in Years) a/b                      2.74                      2.82
2) Project A Project B
Net Present Value $       42,780.71 $       44,835.36
Working:
Present Value of annuity of 1 = (1-(1+1)^-n)/i Where,
= (1-(1+0.19)^-5)/0.19 i 19%
=            3.057635 n 5
Project A Project B
Annual cash flows a $   1,35,000.00 $   1,88,000.00
Present value of
annuity of 1 b            3.057635            3.057635
Present Value of
Annual cash flows c=a*b $   4,12,780.71 $   5,74,835.36
Cost of investment d $   3,70,000.00 $   5,30,000.00
Net Present Value e=c-d $       42,780.71 $       44,835.36
3) Project A Project B
Internal rate of return (IRR) 24.08% 22.74%
Working:
Internal rate of return is the rate at which net present value is zero.
Project A: Project B:
Year Cash flows Year Cash flows
0 $ -3,70,000.00 0 $ -5,30,000.00
1 $   1,35,000.00 1 $   1,88,000.00
2 $   1,35,000.00 2 $   1,88,000.00
3 $   1,35,000.00 3 $   1,88,000.00
4 $   1,35,000.00 4 $   1,88,000.00
5 $   1,35,000.00 5 $   1,88,000.00
IRR 24.08% 22.74%
Project A Project B
4)
Profitability Index                      1.12                      1.08
working:
Project A Project B
Present Value of annual cash flows $   4,12,780.71 $   5,74,835.36
Cost of Investment $   3,70,000.00 $   5,30,000.00
Profitability Index                      1.12                      1.08